Operation Summary for June 6, 2011 - June 8, 2011 |
|
Monday June 6 The day shift began with Operations monitoring the Tevatron (TEV) store 8795 (213.73E30) and the Recycler stash (78.3E10), with the Antiproton Source (Pbar) stacking (30.3E10), with the Main Injector (MI), Linac, and Booster operating normally, with NuMI off for target replacement, with the MiniBooNE and Meson FTBF T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Operators reported, at 8:15 AM, that the MI air pressure was dropping; Mechanical personnel investigated and installed a portable air compressor. Linac personnel rebooted the Linac low-level RF system and a RF station (LRF4). A Linac Klystron RF station (KRF4) tripped off; it reset. During the evening shift, Operations disabled beam to T-978 at 6 PM. Operations terminated store 8795 at 10:07 PM, put the TEV into shot setup, and established store 8796 at 11:42 PM with an average initial luminosity (AIL) of 348.51E30. Operators reported some minor downtimes due to MI-10 vacuum, a BRF4 trip, and MECAR relearning. Tuesday June 6 The midnight shift began with Operations monitoring store 8796 (297.74E30) and the stash (70.0E10), with Pbar stacking (6.5E10), and with MiniBooNE taking beam. During the midnight shift, Operations established beam to T-978 at 4:44 AM. Operators reported that a few MI vacuum incidents contributed to downtime. During the day shift, Operations disabled the T-978 beam and terminated store 8796 at 12:14 PM, put the TEV into shot setup, and established store 8797 at 1:45 PM with an AIL of 336.4E30. Operators prepared Booster and MI for access. Booster personnel removed the NorthStar experiment equipment and came out of access at 3:36 PM. MI remained in access. During the evening shift, MI came out of access at 5:38 PM. Operators conducted a gap clearing kicker investigation that kept MI beam off until about 7 PM. Operators resumed sending beam to MiniBooNE at 6:34 PM and to Pbar at 7:03 PM. Wednesday June 8 The midnight shift began with Operations monitoring store 8797 (88.26E30) and the stash (123.8E10), with Pbar stacking (29.4E10), and with MiniBooNE taking beam. During the midnight shift, at 12:14 AM, the Recycler lost its stash of 123E10 antiprotons following a MI vacuum burst. Operations reported that store 8797 aborted at 7:56 AM when the TEV CAMAC link went down after a Transfer Hall equipment rack lost power. After the MI vacuum recovered, operators transferred antiprotons to the Recycler. Operators prepared the Pelletron for a brief access to retrieve TSP equipment. The Plans for Wednesday FESS personnel have begun repairing the TEV sector A0 cable fault that caused store 8797 to abort. Controls personnel ran some temporary power to the repeater rack; however, two TEV and Switchyard CAMAC crates still need to be restored. The plans for today are uncertain. Operators prepared the TEV for an access. CDF shut down for cryo system repairs and MI is waiting to see if they can go down for six to eight hours for vacuum work. Eventually, the Run Coordinator will want the complex to return to normal stacking, stashing, and storing. On a side note, NuMI hopes to resume normal operations on Saturday morning, 6/11/11. |
|
More Information Comments and Suggestions |
| last modified 06/08/2011 email Fermilab |
| Security, Privacy, Legal | |